The population mean weight of a candy is 10 grams with a standard deviation of 2 grams. A sample of 100 candies is taken. What is the probability that the sample mean is 10.5 grams or lower?
1
Expert's answer
2020-10-23T10:56:03-0400

We need to compute "P(X \le 10.5)". The corresponding z-value needed to be computed are:

"Z = \frac{\bar X- \mu}{\sigma/\sqrt n}= \frac{ 10.5-10}{ 2/\sqrt{100}} = 2.5",

using table, p=0.99379